OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = 4x – 5, what is f(x) ◦ g(x)?

OpenStudy (anonymous):

(2x+3) * (4x-5)

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(4x-5)=2(4x-5)+3\] then algebra

OpenStudy (anonymous):

Is it some symbol in the middle?

OpenStudy (anonymous):

8x-7 or 8x+7 is the answer ?